OpenStudy (anonymous):

A right triangle has legs of length 6 and 11. What is the length of the hypotenuse? PLEASEEEEEEE HELP PRETTY PLEASE

OpenStudy (anonymous):

A^2+B^2=C^2

OpenStudy (anonymous):

The Pythagorean Theorem is:\[a ^{2}+b ^{2}=c ^{2}\]a=6 b=11\[6^{2}=36\]\[11^{2}=121\]\[36+121=157\]\[\sqrt{157}=12.5\]

OpenStudy (mathstudent55):